OpenStudy (anonymous):

how do i find the value of x to the nearest tenths if "tan 34=x/20"

OpenStudy (debbieg):

Solve tan 34=x/20 for x, then punch it up in your calculator. And welcome to Open Study! :)

OpenStudy (anonymous):

how do i solve it

OpenStudy (debbieg):

Rearrange it to isolate the x. Just like if you had, b=x/20 how would you solve that FOR x, eg, rearrange so that you have x=.....?

OpenStudy (debbieg):

tan 34=x/20 What do you need to do to both sides, to have x all alone?

OpenStudy (anonymous):

im confused would I be able to use a calculator if so what do I type in?

OpenStudy (debbieg):

Hold on, one step at a time. Rearrange the equation, so you have x=......

OpenStudy (debbieg):

What if you had the equation: 3=x/20 and you had to solve it for x. what would you do?

OpenStudy (anonymous):

multiply???

OpenStudy (debbieg):

Right! by what?

OpenStudy (debbieg):

Hmmm.... well, you would end up GETTING 60, as your solution for x. But what did you multiply both sides by, to get that?

OpenStudy (anonymous):

3

OpenStudy (debbieg):

no. that would not give you x=60.

OpenStudy (debbieg):

\[\large 3=\frac{ x }{ 20 }\] \[\large 20\cdot 3=\frac{ x }{ 20 }\cdot 20\] \[\large 20\cdot 3=\frac{ x }{ \cancel{20} }\cdot \cancel{20}\] \[\large 60=x\] right?

OpenStudy (debbieg):

Because x/20 is x DIVIDED BY 20. and you want x alone, so you have to "UNDO" that division. so you undo division, by multiplying.

OpenStudy (anonymous):

right

OpenStudy (debbieg):

OK, now instead of: 3=x/20 you have: tan(34) = x/20 so to solve that for x (isolate x) you have to multiply both sides by 20. So you get: x = ??